# Cross-Entropy Games for Language Models: From Implicit Knowledge to General Capability Measures

### v1

@@ 1.1. Explicit Capabilities
For instance, we expect that if we ask for the number of legs of an ant, the answers would be random, but most of them would involve the number 6; and we expect that if we asked the model to produce a single Arabic numeral, it would output the number 6. Hence that number would be part of the explicit knowledge of the LLM.
@@
While many modern LLMs will typically have in their explicit knowledge much general information about how they function internally themselves (e.g. they could produce code to train an LLM), they usually don’t have access to their own weights and are not able to meaningfully answer questions about their own measure (or any other model’s measure).
@@ 1.3.3. LLM Games
Our approach is to embrace the intrinsic subjectivity of this problem by using the same philosophy which leads us to consider the implicit knowledge. This leads us to LLM-based games: situations with LLMs facing LLM-generated contexts and pursuing ascribed objectives, competing or cooperating with one another, with rules enforced by LLM-based arbitration and scores given by LLM measures. As a result of this, optimal play is naturally in the implicit knowledge of the LLM involved (or of the combination of the LLMs involved). The general relevance of LLM-based games to study the implicit knowledge becomes rather natural: such games are about following objectives that result from simple LLM-measurable objectives in an environment created by LLMs, and driven by cooperation and competition with other LLMs.
Games have been used since the inception of AI, lying even at the foundations of the field, with the Turing test (viewed for a long time as a hallmark of AI, as well as a definition) being introduced as an imitation game [Tur50]. In the recent years, games have driven many of the exciting results in the field (see e.g. [MKSLH15, SHSH18, MSBLB17]). However, besides Turing’s foundational works, the choice of the games for AI has traditionally been motivated by socio-historical context (e.g. having been widely played by humans for a long time) rather than intrinsic relevance (unique qualities that make the game worth playing for e.g. other applications); it is hard to argue e.g. that the game of chess is uniquely relevant as a way to achieve general intelligence, or even at achieving any AI objective that does not specifically mention chess. For instance, we are not aware of any perfect-information two-player game (like Go or Chess) designed with general intelligence in mind.
@@
In Section 2, we propose the space of so-called Cross-Entropy Games or Xent Games as an answer to this problem: this space forms a consistent family of games which can be concisely expressed. Informally, Xent Games are about evaluating scenarii that come with ‘scores’ determined by various decorations attached to them and signed cross-entropies evaluations, with signed cross-entropy constraints.
@@
As will be discussed below, the Xent Games can be leveraged to probe the capabilities of LLMs, by exploiting the gap between the explicit and implicit abilities. We rely on the notion of transfer value of a game, to derive a dynamics inspired by evolutionary ideas on the game space, that can be used to probe the general capabilities of LLMs.
@@ 1.4. General Vision and Outline
  • In Section 2, we introduce Xent Games. These are games about strings of tokens evaluated by LLMs, which can be expressed in terms of diagrams, and written down using a simple domain-specific language. We then show that the Xent Game family is the smallest family of games that is stable under a few game-theoretic consistency axioms: in other words, it can be constructed from a single game by iterating a small set of moves.
  • In Section 3, we show that despite being a relatively small family of games, the Xent Game family contains a wealth of examples, in particular all the examples of tasks described in Section 1.2.2. We show that the family contains a number of interesting examples related to those outlined in Section 1.2, and suggest further useful implicit capabilities by LLMs.
@@
  • In Section 5, we discuss the use of the Xent Games as a means to evaluate the abilities of LLMs. The idea is to normalize scores using a base model, and then from a given family of games, build Xent Game Measures: the key idea is that once a scope (a family of Xent Games representing some) has been defined, a minimal covering subfamily can be extracted, leading to a measure.
  • In Section 6, we examine the challenges associated with general capabilities: unbounded scope. Based on game-theoretic considerations and evolution ideas, we propose an algorithm to grow a scope in a systematic and coherent fashion. This leads in particular to a theoretically-motivated path to measure the general capabilities of LLMs.
@@ Acknowledgements
Many insights presented in this paper have emerged from works done by the first author in (past and ongoing) collaborations with Diego Dorn, Franck Gabriel, Vassilis Papadopoulos, Arthur Renard, Marco Tuccio, and Jérémie Wenger on Large Language Models, with whom key ideas were discussed and investigated.
In addition, the authors would like to thank Emmanuel Abbé, Apoorv Agarwal, Alberto Bietti, Gloria Capano, Tarun Chitra, Jordan Cotler, Mario Geiger, Nicola Greco, Bara Hudcová, Arthur Jacot, Niels Linnemann, João Penedones, and Stanislav Smirnov for interesting discussions, as well as the participants to the Quine seminar and Demeco workshop for insightful questions and remarks (in particular Clément Moulin-Frier).
@@ 2. Xent Games
In Section 1.3.3, the idea to use games to elicit the implicit measure of LLMs was introduced, leading the question of finding a suitable space of games (Problem 10). In this section, we introduce Cross-Entropy Games, or Xent Games as an answer to this question.
@@ 2.1. Informal Description and Goals
Xent Games are single- or multiplayer turn-by-turn (with a finite number of turns) text-based games involving a reference LLM measure $\mathcal{M}$ to assign, from the cross-entropy action $\mathcal{S}_{\mathcal{M}}$:
@@
Xent Games are general-sum imperfect information games with complete information (the rules are known to all players), though many interesting examples can already be found among perfect information games, in particular the single-player and the two-player zero-sum ones (see Section 3).
@@ 2.1.1. Nature of Xent Games
Informally, one could say that Xent Games are about studying scenarii (in the form of combinations of strings produced by the players), and weighing their plausibility in terms of cross-entropy measures; in other words, playing Xent Games is tantamount to identifying plausible paths in spaces of scenarii that fulfill constraints that are either explicit a priori, or that come from other players’ moves (which themselves can be produced by a perfect information or an imperfect information setup). Roughly speaking, these can be described as the ‘games that live in the minds of LLMs’.
As a result, Xent Games are really about path-finding in a complex string environment dictated by cross-entropy measures (via the ‘judge model’, see 2.2.2 below), with the idea that such environments are partly static, partly dynamical, being influenced by other agents; the scenarii are formed by concatenation and cuts of strings, which are either given a priori, generated by a ‘map-generator’ (which we sometimes refer to as ‘the stories’), or by the various players. This informal description should make it unsurprising that the class of Xent Games is in fact very rich (see e.g. Section 3 for a few examples).
@@
It should be noted that for long-range natural language cases, some care must be taken to ensure games are nontrivial: minimizing the cross-entropy of a continuation (the so-called ‘MAP-continuation’) to a given text with no constraints can lead to trivial repetitive outputs [StBy19, MCV20]. While it is important to have these considerations in mind, it is important to note that in many cases (particularly reasoning), minimal cross-entropy solutions are in fact useful and non-degenerate [SWLL24, SYCYL24]. Finally, and most importantly, these issues are not critical here: what matters for the purpose of answering Problem 10 is to identify ‘enough’ games.
@@ 2.1.2. Xent Game Space
Because of their common structure Xent Games are tightly related to one another: via simple operations on a Xent Game, one can end up with a different one; in fact, the minimality of the Xent Game Space under a few basic game-theoretic operations essentially means one can go from any Xent Game to any other one by a sequence of natural ‘moves’ in the game space (see Section 2.3).
The Xent Games are designed to be played by LLMs; typically when using a Xent Game to evaluate or train a model, a ‘main character player’ will be picked among the other NPC players (following the videogame terminology for non-playable characters). Each Xent Game instance (i.e. ‘New Game’, in videogame terminology) is typically built upon a sampled string ‘context’ (i.e. the ‘Game Map’), which can be generated by an LLM with a prompt.
In the subsequent Sections, Xent Games will be used as a means to assess and develop the abilities of an LLM playing them. In particular, measures of the playability value (see Section 4.2) and of the transfer value (see Section 4.3) will be introduced, that will allow us to explore the space of Xent Games in a principled way.
@@ 2.2. Xent Games: Context, Moves, Constraints, and Rewards
As discussed above, Xent Games are turn-by-turn $n$-player (with $n\geq 1$) complete information games (the rules are known to everyone) played in the space of strings (i.e. they are text-based), running for a fixed finite number of steps. In this subsection, we define Xent Games mathematically. The programming language and the graphical representation used to build Xent Games are presented in Sections 2.4 and 2.5 respectively; note that for practical reasons, the language and representation only implement a subset of the Xent Games (with a limit on the number of players, of turns, etc.) and that, to make the writing of relevant games easier, they include some syntactic sugar and extra functions (which is not discussed here).
@@ 2.2.1. Game Metadata
  • The specification of the ‘judge’ model $\mathcal{J}$ (see 2.2.2) used for the rewards and the constraint enforcement.
@@
  • The specification of string constants used as part of the context and as part of prompt to generate the context.
  • The specification of size constants used as constraints on the sizes of the player’s moves.
@@
In addition to the above, each game comes with a special piece of metadata, which is the session seed, used for the model generation of the metadata. This is typically set externally at the time of running of the games.
@@ 2.2.2. Judge Model
This doesn’t mean that playing with relatively weak judge models is absolutely uninteresting: finding a way around naive enforcement of rules can actually be an interesting task. In fact, an interesting choice for a judge model can be an ‘umbrella model’, that hosts several models of various capabilities under the hood, with possibly an activation prefix token to specify if an unusual model is to be used. These kinds of constructions can be useful for e.g. contrastive problems (see Section 1.2.2 above), where one tries e.g. to maximize the plausibility for a strong model while reducing the plausibility for a weak model.
For the sake of concreteness and simplicity, the model $\mathcal{J}$ can be assumed to be a strong pre-trained LLM throughout this paper. Still, as discussed in Section 7.2.1, the role of $\mathcal{J}$ should not be under-estimated.
@@ 2.2.3. String Space and Operations
The Xent Games are inherently text based: the players process the rules (which are written in code), receive text information and play text moves. The string space at any time of the game is a fixed set of variables, which are all strings, initialized by default to be the empty string. The set of variables includes the context, the moves of the players and all intermediate variables relevant. Outside of initialization and player moves, the string space can evolve (via assignment to existing variables from existing variable), via the following simple ‘cat and cut’ operations:
  • Cat: this is simply the concatenation of two strings (at the token level), typically denoted by $s1+s2$; in practice, note that this means that a space character is sandwiched between $s1$ and $s2$.
  • Cut: this corresponds to the ‘head’ and ‘tail’ of the ‘split’ operation, i.e. $s//t$ denotes the substring of $s$ coming before the first occurrence of $t$ (not including $t$; corresponding to $s$ if $t$ does not appear in $s$ or if $t$) and $s\%t$ denotes the substring of $s$ coming after the first occurrence of $t$ (not including $t$; corresponding to the empty string $t$ does not appear in $s$). By convention, we say that $t$ does not appear in $s$ if it is empty.
It is easy to see that this space of allowed operations can in particular be used to do matching and replacement for basic patterns (for a fixed maximal number of appearances of the pattern).
@@ 2.2.5. Moves and Constraints
    • A set of variables of the string space.
@@ 2.2.6. Game Operations
Assuming the metadata being specified, the game operations go as follows:
@@
  • The game loop runs a fixed sequence of basic steps (determined in advance, and of length not exceeding the maximal number specified in the metadata). Each basic step consists in one of the following:
@@ 2.2.7. Game Rules and Code
As will be explained in Section 2.4 below, the Xent Games can be written using a domain-specific programming language and a graphical language.
@@ 2.2.8. Xent Game Space: Key Design Philosophy
  • These are games where agents play with the implicit measure $\mathcal{S}_{\mathcal{J}}$ of a given model, trying to accommodate constraints dictated by it and to optimize information-theoretic scores related to it, and thinking strategically about the other players (for multi-player games). In a sense, they are games of string composition with fixed, cooperative, and competitive constraints, where the basic tools for composing strings are cat/cut operations.
  • While the games revolve around optimizing quantities based cross-entropy, playing them does not need to explicitly output cross-entropy or probability numbers in text, which is not very natural for text-based outputs.
  • The design is so that games run for a fixed amount of time (provided that feasible moves can be found), and that their syntactic validity just corresponds to the individual independent validity of each step, making it easy to ‘mix’ games, by step combination.
@@
  • Though clearly limiting, the use of signed xent sums rather than more general functions of xents (or even of arbitrary linear combinations) makes normalization questions easier (see 5.2 below), and it alleviates an over-reliance on subtle arithmetic operations which do not correspond to conceptually interesting implicit measure questions. Furthermore, the signed xent sums are definitely close in spirit to quantities that would naturally lend themselves to information-theoretic interpretations (see e.g. the ‘bits back’ coding introduced in [HiVC93]).
  • It should not be expected that all Xent Games bring transfer value, or even that they are playable. A key (heuristic) thesis of this paper though is that the space of Xent Games is so rich that it is in fact relatively easy to find ‘good’ games in various senses that we discuss in Sections 4.2, 6, and 7 below.

### v2

@@ 1.1. Explicit Capabilities
For instance, we expect that if we ask for the number of legs of an ant, the answers would be random, but most of them would involve the number 6; and we expect that if we asked the model to produce a single Arabic numeral, it would output the number 6. Hence the fact that ants have 6 legs would be part of the explicit knowledge of the LLM.
@@
While many modern LLMs will have in their explicit knowledge general information about how they function internally (e.g. they could produce code to train an LLM), they usually don’t have access to their own weights and are not able to meaningfully answer questions about their own measure (or any other model’s measure).
@@ 1.3.3. LLM Games
Our approach is to embrace the intrinsic subjectivity of this problem by using the same philosophy which lead us to consider implicit knowledge. This leads us to LLM-based games: situations with LLMs facing LLM-generated contexts and pursuing ascribed objectives, competing or cooperating with one another, with rules enforced by LLM-based arbitration and scores given by LLM measures. As a result of this, optimal play is naturally in the implicit knowledge of the LLM involved (or of the combination of the LLMs involved). The general relevance of LLM-based games to study the implicit knowledge of LLMs becomes rather natural: such games are about following objectives that result from simple LLM-measurable scoring in an environment created by LLMs, and driven by cooperation and competition with other LLMs.
Games have been used since the inception of AI, lying at the foundations of the field, with the Turing test (viewed for a long time as a hallmark of AI, as well as a definition) being introduced as an imitation game [Tur50]. In the recent years, games have driven many of the exciting results in the field (see e.g. [MKSLH15, SHSH18, MSBLB17]). However, besides Turing’s foundational works, the choice of the games for AI has traditionally been motivated by socio-historical context (e.g. having been widely played by humans for a long time) rather than intrinsic relevance (unique qualities that make the game worth playing); it is hard to argue e.g. that the game of chess is uniquely relevant as a way to achieve general intelligence, or even at achieving any AI objective that does not specifically mention chess. For instance, we are not aware of any perfect-information two-player game (like Go or Chess) designed with general intelligence in mind.
@@
In Section 2, we propose the space of so-called Cross-Entropy Games or Xent Games as an answer to this problem: this space forms a consistent family of games which can be concisely expressed. Informally, Xent Games are about evaluating scenarii that come with ‘scores’ determined by signed cross-entropies evaluations with signed cross-entropy constraints.
@@
As will be discussed below, the Xent Games can be leveraged to probe the abilities of LLMs by exploiting the gap between the explicit and implicit capabilities. We rely on the notion of transfer value of a game, to derive a dynamics inspired by evolutionary ideas on the game space, that can be used to probe the general capabilities of LLMs.
@@ 1.4. General Vision and Outline
  • In Section 2, we introduce Xent Games. These are games about strings of tokens evaluated by LLMs, which can be expressed in terms of diagrams, and written down using a simple domain-specific language. We then show that the Xent Game family is the smallest family of games that is stable under a few game-theoretic consistency axioms: in other words, the Xent Game family can be constructed from a single game by iterating a small set of moves.
  • In Section 3, we show that despite being a relatively small family of games, the Xent Game family contains a wealth of examples, in particular all the examples of tasks described in Section 1.2.2. We show that the family contains a number of interesting examples related to those outlined in Section 1.2, and suggest further useful implicit capabilities of LLMs.
@@
  • In Section 5, we discuss the use of the Xent Games as a means to evaluate the abilities of LLMs. The idea is to normalize scores using a base model, and then from a given family of games, build Xent Game Measures: the key idea is that once a scope (a family of Xent Games representing some abilities) has been defined, a minimal covering subfamily can be extracted, leading to the creation of a measure.
  • In Section 6, we examine the key challenge associated with general capabilities: unbounded scope. Based on game-theoretic considerations and evolutionary ideas, we propose an algorithm to grow a scope in a systematic and coherent fashion. This leads, in particular, to a theoretically-motivated path to measure the general capabilities of LLMs.
@@ Acknowledgements
Many insights presented in this paper have emerged from work done by the first author in (past and ongoing) collaborations with Diego Dorn, Franck Gabriel, Vassilis Papadopoulos, Arthur Renard, Marco Tuccio, and Jérémie Wenger on Large Language Models, with whom key ideas were discussed and investigated.
In addition, the authors would like to thank Emmanuel Abbé, Apoorv Agarwal, Alberto Bietti, Gloria Capano, Tarun Chitra, Jordan Cotler, Mario Geiger, Nicola Greco, Leonard Hardiman, Bara Hudcová, Arthur Jacot, Niels Linnemann, João Penedones, and Stanislav Smirnov for interesting discussions and comments on earlier version of the manuscript, as well as the participants to the Quine seminar and Demeco workshop for insightful questions and remarks (in particular Clément Moulin-Frier).
@@ 2. Xent Games
In Section 1.3.3, the idea to use games to elicit the implicit measure of LLMs was introduced, leading to the problem of finding a suitable space of games (Problem 10). In this section, we introduce Cross-Entropy Games, or Xent Games as an answer to this question.
@@ 2.1. Informal Description and Goals
Xent Games are single- or multi-player turn-by-turn (with a finite number of turns) text-based games involving a reference LLM measure $\mathcal{M}$ to assign, from the cross-entropy action $\mathcal{S}_{\mathcal{M}}$:
@@
Xent Games are general-sum imperfect information games with complete information (the rules are known to all players), though many interesting examples can already be found among perfect information games, in particular single-player and two-player zero-sum games (see Section 3).
@@ 2.1.1. Nature of Xent Games
Informally, one could say that Xent Games are about studying scenarii (in the form of combinations of strings produced by the players), and weighing their plausibility in terms of cross-entropy measures; in other words, playing Xent Games is tantamount to identifying plausible paths in spaces of scenarii that fulfill constraints that are either explicit a priori, or that come from other players’ moves (which themselves can be produced by a perfect information or an imperfect information setup). Roughly speaking, these can be described as ‘games that live in the minds of LLMs’.
As a result, Xent Games are really about path-finding in a complex string environment dictated by cross-entropy measures (via the ‘judge model’, see 2.2.2 below), with the idea that such environments are partially static and partially dynamic due to the influence of other agents. The scenarii are formed by concatenation and cuts of strings, which are either given a priori, generated by a ‘map-generator’ (which we sometimes refer to as ‘the stories’), or elicited from the various players. This informal description should make it unsurprising that the class of Xent Games is in fact very rich (see e.g. Section 3 for a few examples).
@@
It should be noted that for long-range natural language cases, some care must be taken to ensure games are nontrivial: minimizing the cross-entropy of a continuation (the so-called ‘MAP-continuation’) of a given text with no constraints can lead to trivial repetitive outputs [StBy19, MCV20]. While it is important to have these considerations in mind, it should be noted that in many cases (particularly reasoning), minimal cross-entropy solutions are in fact useful and non-degenerate [SWLL24, SYCYL24]. Finally, and most importantly, these issues are not critical here: what matters for the purpose of answering Problem 10 is just to identify ‘enough’ games.
@@ 2.1.2. Xent Game Space
Because of their common structure, Xent Games are tightly related to one another: via simple operations on a Xent Game, one can end up with a different one; in fact, the minimality of the Xent Game Space under a few basic game-theoretic operations essentially means one can go from any Xent Game to any other one by a sequence of natural ‘moves’ in the game space (see Section 2.3).
The Xent Games are designed to be played by LLMs; typically when using a Xent Game to evaluate or train a model, a ‘main character player’ will be picked from among the other NPC players (following the videogame terminology for non-playable characters). Each Xent Game instance (i.e. ‘New Game’, in videogame terminology) is typically built upon a sampled string ‘context’ (i.e. the ‘Game Map’), which can be generated by an LLM with a prompt.
In the subsequent Sections, Xent Games will be used as a means to assess and develop the abilities of an LLM playing them. In particular, measures of the playability value (see Section 4.2) and of the transfer value (see Section 4.3) will be introduced, which will allow us to explore the space of Xent Games in a principled way.
@@ 2.2. Xent Games: Context, Moves, Constraints, and Rewards
As discussed above, Xent Games are turn-by-turn $n$-player (with $n\geq 1$) complete information games (the rules are known to everyone) played in the space of strings (i.e. they are text-based), running for a fixed finite number of steps. In this subsection, we define Xent Games mathematically. The programming language and the graphical representation used to build Xent Games are presented in Sections 2.4 and 2.5 respectively; note that for practical reasons, the language and representation only implement a subset of the Xent Games (with a limit on the number of players, of turns, etc.) and that, to make the writing of relevant games easier, they include some syntactic sugar and extra functions.
@@ 2.2.1. Game Metadata
  • The specification of the ‘judge’ model $\mathcal{J}$ (see 2.2.2) used for rewards and constraint enforcement.
@@
  • The specification of string constants used as part of the context and as part of the prompts to generate the context.
  • The specification of size constants used as constraints on the sizes of the players’ moves.
@@
In addition to the above, each game comes with a special piece of metadata, which is the session seed, used for the model generation of the context. This is typically set externally from the rest of the metadata.
@@ 2.2.2. Judge Model
This doesn’t mean that playing with relatively weak judge models is absolutely uninteresting: finding a way around naive enforcement of rules can actually be an interesting task. In fact, an interesting choice for a judge model can be an ‘umbrella model’ which hosts several models of various capabilities under the hood, with possibly an activation prefix token to specify if a particular model is to be used. These kinds of constructions can be useful for e.g. contrastive problems (see Section 1.2.2 above), where one may try to maximize plausibility for a strong model while reducing plausibility for a weak model.
For the sake of concreteness and simplicity, the model $\mathcal{J}$ can be assumed to be a strong, pre-trained LLM throughout this paper. Still, as discussed in Section 7.2.1, the role of $\mathcal{J}$ should not be under-estimated.
@@ 2.2.3. String Space and Operations
The Xent Games are inherently text based: the players process the rules (which are written in code), receive text information, and play text moves. The string space at any time of the game is a fixed set of variables, which are all strings, initialized by default to be the empty string. The set of variables includes the context, the moves of the players and all relevant intermediate variables. Outside of initialization and player moves, the string space can evolve (due to assignment to existing variables from existing variable), via the following simple ‘cat’ and ‘cut’ operations:
  • Cat: this is simply the concatenation of two strings (at the token level), denoted by $s1+s2$. Note that, in practice, this means that a space character is sandwiched between $s1$ and $s2$.
  • Cut: this corresponds to the ‘head’ and ‘tail’ of a ‘split’ operation. $s//t$ denotes the substring of $s$ coming before the first occurrence of $t$ (not including $t$; corresponding to $s$ if $t$ does not appear in $s$) and $s\%t$ denotes the substring of $s$ coming after the first occurrence of $t$ (not including $t$; corresponding to the empty string $t$ does not appear in $s$). By convention, we say that $t$ does not appear in $s$ if it is empty.
It is easy to see that this space of allowed operations can be used to do matching and replacement for basic patterns (for a fixed maximal number of appearances of the pattern).
@@ 2.2.5. Moves and Constraints
    • A set of variables in the string space.
@@ 2.2.6. Game Operations
Once the metadata is specified, the game operations go as follows:
@@
  • The game loop runs a fixed sequence of basic steps (determined in advance, and of length not exceeding the maximal number specified in the metadata). Each basic step consists of one of the following:
@@ 2.2.7. Game Rules and Code
As will be explained in Section 2.4 below, Xent Games can be written using a domain-specific programming language and a graphical language.
@@ 2.2.8. Xent Game Space: Key Design Philosophy
  • These are games where agents play with the implicit measure $\mathcal{S}_{\mathcal{J}}$ of a given model, trying to accommodate constraints dictated by it and to optimize information-theoretic scores related to it, while thinking strategically about the other players (for multi-player games). In a sense, they are games of string composition with fixed, cooperative, and competitive constraints, where the basic tools for composing strings are cat/cut operations.
  • While the games revolve around optimizing cross-entropy based quantities, playing them does not need to explicitly output cross-entropy or probability numbers in text, which would not be very natural for text-based outputs.
  • Games are designed such that they run for a fixed amount of time (provided that feasible moves can be found), and their syntactic validity simply corresponds to the individual independent validity of each step, making it easy to ‘mix’ games by step combination.
@@
  • Though clearly limiting, the use of signed xent sums rather than more general functions of xents (or even of arbitrary linear combinations) makes normalization questions easier (see 5.2 below), and it alleviates an over-reliance on subtle arithmetic operations which do not correspond to conceptually interesting implicit measure questions. Furthermore, signed xent sums are definitely close in spirit to quantities that would naturally lend themselves to information-theoretic interpretations (see e.g. the ‘bits back’ coding introduced in [HiVC93]).
  • It should not be expected that all Xent Games bring transfer value, or even that they are playable. A key (heuristic) thesis of this paper is that the space of Xent Games is so rich that it is in fact relatively easy to find ‘good’ games in various senses that we discuss in Sections 4.2, 6, and 7 below.